# Building Trustworthy AI by Addressing its 16+2 Desiderata with Goal-Directed Commonsense Reasoning

### v1

@@ 3 Addressing 16+2 desiderata with s(CASP)
Commonsense reasoning systems, including s(CASP), represent knowledge as predicates. To communicate with humans, these predicates must be translated to a natural language back and forth. Machine translation of languages has been shown to be done well with pattern matching. Thus LLMs are able to perform this task with high accuracy. The reliable chatbots referenced in Desiderata LABEL:des:06 illustrates this very well. These reliable chatbot applications (?) usually target some specific tasks, and achieve the goal by collecting as much necessary information as they can from the human users under the guidance of s(CASP)’s knowledge base. Take the restaurant recommendation chatbot as an example. A customer’s food preferences, budget, family-friendliness requirements, etc., must be known before a recommendation is given. This is encoded as part of s(CASP)’s knowledge. The chatbot uses an LLM to translate user dialogs to gather these requirements represented as predicates, then makes a recommendation using its encoded logic. It is also able to explain (in a natural language) why it made a specific recommendation. By using an LLM as a semantic parser, the s(CASP)-based chatbot is able to have linguistic proficiency similar to a human.
@@
Discussion: As mentioned in Desiderata LABEL:des:06, while LLMs are very adaptable in their manner of speech they are also prone to being jailbroken. Additionally, because of the quantity of data LLMs have consumed, they often need to be fine-tuned to cover a specific domain. Cyc can cover a variety of domains and does have some functionality for natural language, but it is noted to be rudimentary. In our opinion, the best approach to making AI lingual is to leverage LLMs.
Table 1: Desiderata accomplishment by Cyc, Deep-Learning (DL), and s(CASP). Including s(CASP) features.
Cyc DL s(CASP)
Feature / Tool
LABEL:des:01 Explanation Justification Trees
LABEL:des:02 Deduction ! Base s(CASP)
LABEL:des:03 Induction ! ! Default Rules + FOLD
LABEL:des:04 Analogy ! ! ! Base s(CASP)
LABEL:des:05 Abductive Reasoning Abducibles
LABEL:des:06 Theory of Mind Chatbot
LABEL:des:07 Quantifier-fluency Goal-Directed Execution
LABEL:des:08 Modal-fluency Deontic Logic
LABEL:des:09 Defeasibility Default Rules + FOLD
LABEL:des:10 Pro and Con Arguments ASP Multiple Worlds
LABEL:des:11 Contexts VECSR
LABEL:des:12 Meta-knowledge and Meta-reasoning Justification Trees
LABEL:des:13 Explicitly Ethical ! Constraints
LABEL:des:14 Sufficient Speed ! VECSR
LABEL:des:15 Sufficiently Lingual and Embodied Chatbot
LABEL:des:16 Broadly and Deeply Knowledgable ! FOLD, Chatbot, VECSR
LABEL:des:17 Inconsistency Detection ! Constraints
LABEL:des:18 Multiple Possible Worlds ! ASP Multiple Worlds
[…]
@@
An effective AI must possess or access a broad and deep foundation of world knowledge, from commonsense to specialized domains, enabling it to interact meaningfully and contextually with people. While memorizing facts is less critical in the internet age, a trustworthy AI must skillfully retrieve, understand, and reason about information in real time, drawing on structured sources and web services. Importantly, it should interpret the meaning, trustworthiness, and implications of knowledge, using logical, analogical, and inferential reasoning as humans do.
s(CASP)
@@
There are multiple ways to incorporate new information into a s(CASP) system. The FOLD algorithms can make inductions from multiple forms of data, including traditional datasets and even images (?). NeSyFOLD in particular demonstrates how the advantages of deep learning (such as robust image analysis) can be leveraged with s(CASP) to add explainability and transparent decision analysis to AI systems. The ability to learn from data is critical to build a system that is broadly and deeply knowledgeable, as there is too much information available in the connected modern age to ever encode by hand. Data can also be added from an LLM as in the chatbot research or from simulated environments as in VECSR and reasoned over using the s(CASP) system.
@@
Discussion: LLMs consume vast quantities of data, however there is growing concern that LLMs will run out of data before achieving true reasoning (?). Because of the large quantity of data LLMs must be trained on, there can often be incorrect or low-quality data included, which is then reflected in wrong answers provided by the LLM. Cyc’s solution to this has historically been the creation of its bespoke, human-groomed knowledge base, which trades man-hours for accuracy. However, it is not possible to encode the breadth of knowledge LLMs gain through training by hand in an amount of time that allows for it to be kept up to date with the changing world. Some compromise must be made to ensure reasonable accuracy of the data within a reasonable amount of time.
[…]
@@
An effective AI system must be able to detect inconsistencies and enforce invariants, something that is easy for humans. The AI system, for example, should be able to tell that a person cannot sit and stand at the same time (inconsistency) or that an alive human must always breathe (invariant).
s(CASP)
@@
Inconsistencies and invariants can be modeled quite simply as global constraints in ASP, and are supported in s(CASP). Examples above will be coded as:
false :- person(X), sit(X), stand(X).
@@
Discussion: The ability to specify inconsistencies has not been presented as a desideratum by ? (?). However, it is an extensively used feature in commonsense reasoning. Humans learn a vast number of such inconsistencies and invariants and use them to derive conclusions. LLMs do not have an explicit representation of inconsistencies, and the lack of enforcing consistency is a source of hallucinations. Cyc’s reasoners have to have inconsistency detection embedded, and it is likely that the various reasoners incorporate it.
[…]
An effective AI system should reason about possible worlds that can simultaneously exist. Humans have the ability to simultaneously represent multiple possible worlds in their mind and reason over each separately. For example, the real world and the world of cartoons are two different worlds with parts that are common (fish can swim in both worlds, for example) and certain parts that are inconsistent (fish talk like humans in one and not the other).
s(CASP)
In s(CASP), multiple possible worlds are easily represented through even loops over negation as discussed in the Abduction and Context desiderata. s(CASP) explicitly supports co-inductive reasoning (which relies on greatest fixpoint semantics), essential to representing multiple possible worlds.
Discussion: LLMs do not have an explicit notion of multiple possible worlds, and this again can be a source of hallucinations. In fact, machine learning systems are often not even aware of such nuances, and conflating of multiple worlds can be a source of inaccuracy and hallucination. Cyc supports contexts which can be thought of as different worlds, however, additional logic is needed to distinguish between different contexts and whether one context is a sub-context of another one. Lack of support for multiple possible worlds is a legacy of Russell’s insistence on only relying on well-founded constructs (?). ASP/s(CASP) break away from this dogma and support non-well-founded semantics, crucial to supporting commonsense reasoning.
@@ 4 Conclusion
Table 1 shows that the 16 (+2) desiderata that trustworthy AI systems must support are almost all entailed by s(CASP), by exploiting its basis or other tools/features developed recently. While ASP/s(CASP) can be thought of as modeling commonsense reasoning in a formal/structured manner, a solution like Cyc relies on its large human-created knowledge base and proprietary reasoners. Moreover, s(CASP) has been tested in combination with other tools, such as ILP, LLM, and simulators, achieving results comparable to those obtained by state-of-the-art approaches.
As future work, we believe that in order to consider s(CASP) a truly reliable AI it is necessary to improve its execution efficiency, applying techniques based on static analysis such as the dynamic consistency check by ? (?), and to build a large foundation of commonsense knowledge like Cyc already possesses. Nevertheless, s(CASP) has already proven its usefulness in specific applications, such as the work done by Forsante Corp of Finland on the Clozapine drug delivery guide, which has been certified under the EU Medical Device Regulation (class 2b) (?). Thus s(CASP) is used commercially while still being open source and freely available for anyone looking to create trustworthy AI applications.
Acknowledgments

### v2

@@ 1 Multiple Possible Worlds ! ASP Multiple Worlds[6]
s(CASP): Commonsense reasoning systems, including s(CASP), represent knowledge as predicates. To communicate with humans, these predicates must be translated to and from natural language. Machine translation of languages has been shown to be done well with pattern matching. Thus, we employ LLMs to perform this task with high accuracy. The reliable chatbots referenced in Desideratum 1 illustrate this very well. These reliable chatbot applications (Zeng et al., 2023) usually target some specific tasks, and achieve the goal by collecting as much necessary information as they can from the human users under the guidance of s(CASP)’s knowledge base. Take the restaurant recommendation chatbot (Zeng et al. (2023) as an example. A customer’s food preferences, budget, family-friendliness requirements, etc., must be known before a restaurant recommendation is given. This is encoded as part of s(CASP)’s knowledge. The chatbot uses an LLM to translate user dialogs to gather these requirements represented as predicates, then makes a recommendation using its encoded logic. It is also able to explain in natural language why it made a specific recommendation. By using an LLM as a semantic parser, the s(CASP)-based chatbot is able to have linguistic proficiency similar to a human.
Additionally, s(CASP) has been used in a fully embodied way by the VECSR system described in Desideratum 1. Using a s(CASP)-controlled agent in an simulation environment, s(CASP) can be used to reason over a wide variety of external stimuli and perform tasks in an embodied way. For complete details, see Tudor et al. (2025).
@@
Discussion: As mentioned in Desideratum 1, while LLMs are very adaptable in their manner of speech they are also prone to being jail-broken. Additionally, because of the quantity of data LLMs have consumed, they often need to be fine-tuned to cover a specific domain. Cyc can cover a variety of domains and does have some functionality for natural language, but it is noted to be rudimentary. In our opinion, the best approach to making AI lingual is to leverage LLMs as translation devices for converting text into a logic-based formalism and vice versa.
@@
#16 Broadly and Deeply Knowledgeable: An effective AI must possess or access a broad and deep foundation of world knowledge, from commonsense to specialized domains, enabling it to interact meaningfully and contextually with people. While memorizing facts is less critical in the internet age, a trustworthy AI must skillfully retrieve, understand, and reason about information in real time, drawing on structured sources and web services. Importantly, it should interpret the meaning, trustworthiness, and implications of knowledge, using logical, analogical, and inferential reasoning as humans do.
@@
s(CASP): There are multiple ways to incorporate new information into a s(CASP) system. The FOLD algorithms can induce from multiple forms of data, including traditional datasets and even images (Padalkar et al., 2024). NeSyFOLD in particular demonstrates how the advantages of deep learning (such as robust image analysis) can be leveraged with s(CASP) to add explainability and transparent decision analysis to AI systems. The ability to learn from data is critical to build a system that is broadly and deeply knowledgeable, as there is too much information available in the connected modern age to ever encode by hand. Data can also be added from an LLM as in the reliable chatbot research by Zeng et al. (2024) or from simulated environments as in VECSR and reasoned over using the s(CASP) system.
@@
Discussion: LLMs consume vast quantities of data, however there is a growing concern that LLMs will run out of data before achieving true reasoning (Villalobos et al., 2022). Because of the large quantity of data LLMs must be trained on, there can often be incorrect or low-quality data included, which is then reflected in wrong answers provided by the LLM. Cyc’s solution to this has historically been the creation of its bespoke, human-groomed knowledge base, which trades man-hours for accuracy. However, it is not possible to encode the breadth of knowledge LLMs gain through training by hand in an amount of time that allows for it to be kept up to date with the changing world.
@@
#17 Inconsistency Detection: An effective AI system must be able to detect inconsistencies and enforce invariants, something that is easy for humans. The AI system, for example, should be able to tell that a person cannot sit and stand at the same time (inconsistency) or that an alive human must always breathe (invariant).
@@
s(CASP): Inconsistencies and invariants can be modeled quite simply as global constraints in ASP, and are supported in s(CASP). Examples above will be coded as:
@@
Discussion: The ability to specify inconsistencies has not been presented as a desideratum by Lenat & Marcus (2023). However, it is an extensively used feature in commonsense reasoning. Humans learn a vast number of such inconsistencies and invariants and use them in reasoning. LLMs do not have an explicit representation of inconsistencies, and the lack of enforcing consistency is a source of hallucinations. Cyc’s reasoners have to have inconsistency detection embedded, and it is likely that the various reasoners incorporate it.
#18 Multiple Possible Worlds: An effective AI system should reason about possible worlds that can simultaneously exist. Humans have the ability to simultaneously represent multiple possible worlds in their mind and reason over each separately. For example, the real world and the world of cartoons are two different worlds with parts that are common (fish can swim in both worlds, for example) and certain parts that are inconsistent (fish talk like humans in the cartoon world and not in the real world).
s(CASP): In s(CASP), multiple possible worlds are easily represented through even loops over negation as discussed in the Abduction and Context desiderata. s(CASP) explicitly supports co-inductive reasoning (which relies on greatest fixpoint semantics), essential to representing multiple possible worlds.
Discussion: LLMs do not have an explicit notion of multiple possible worlds, and this again can be a source of hallucinations. In fact, machine learning systems are often not even aware of such nuances, and conflating of multiple worlds can be a source of inaccuracy and hallucination. Cyc supports contexts which can be thought of as different worlds, however, additional logic is needed to distinguish between different contexts and to indicate whether one context is a sub-context of another one. Lack of support for multiple possible worlds is a legacy of Russell’s insistence on only relying on well-founded constructs (Gupta et al., 2022). ASP/s(CASP) break away from this dogma and support non-well-founded semantics, crucial to supporting commonsense reasoning.
@@ 4 Conclusion
Table 1 shows that the 16 (+2) desiderata that trustworthy AI systems must support could be supported by s(CASP) either directly or through recent applications. As future work, we believe that in order for s(CASP) to enable truly reliable AI it is necessary to improve its execution efficiency, applying techniques such as the dynamic consistency check by Arias et al. (2022b), and to build a large foundation of commonsense knowledge like Cyc already possesses. Brittleness is another issue in symbolic knowledge representation, i.e., KB construction is error prone, and a small change can drastically change the meaning of a KB. This is somewhat alleviated by providing explanations and program debugging/tracing tools created for s(CASP). However, what will further help are tools that can statically analyze, for example, if adding a constraint will make a KB inconsistent. Nevertheless, s(CASP) has already proven its usefulness in specific applications, such as the IRMA system by Forsante Corp of Finland for automating the Clozapine drug delivery guidelines. IRMA is the first logic-based system certified under the EU Medical Device Regulation (class 2b) (Forsante Oy, 2025). Thus s(CASP) is used commercially while still being open source and freely available for anyone looking to create trustworthy AI applications.
Acknowledgements